Beam status
MCR News
21 February 2026 13:31
We briefly established base rate beam before being tripped back off by the
TS2 decoupled moderator. Target staff are on the way to site to
investigate. We will run beam to TS1 only in the meantime.
21 February 2026 11:49
The water leak in the outer synchotron has been fixed. Any associated
equipment is back on and working. We will lock the area back up and run
the main magnet power supply back up to full level. Update at 13:00.
21 February 2026 09:13
We have found a water leak from a magnet insulator in the synchrotron.
This was causing the problem we had earlier with the beam loss monitors.
Mechanical staff are on site to change the insulator. Update at 11:30.
21 February 2026 07:10
Beam is currently off whilst our diagnostics specialists address an issue
the synchrotron beam loss monitors.
20 February 2026 21:22
Attention all users:
Beam to both target stations will be intermittent over the next few hours
to allow our Physicists to set up beam to Target Station 2.
20 February 2026 21:16
Attention TS2 users:
Shutter checks are now complete.
The following beamlines are cleared for beam:
W6 LET
W7 NNIMROD
W1 CHIPIR NOT CLEARED-REQUIRES A RADIATION SURVEY
W5 IMAT NOT CLEARED-REQUIRES A RADIATION SURVEY
E1 ZOOM (SHUTTER IS OPEN)
E2 SANS2D
E3 POLREF
E4 INTER (SHUTTER IS OPEN)
E5 OFFSPEC
E6 LARMOR
E8 WISH (SHUTTER IS OPEN)
All cleared beamlines have control of the shutters.
20 February 2026 15:43
We are shortly coming to the end of the TS2 machine checks. Full beam will
be reinstated to TS1 when they are finished. At 20:00 hrs this evening TS2
beam alignment will commence with base rate beam. Beam will be
intermittent to both targets whilst this is conducted. Any problems,
please ring the MCR.
20 February 2026 14:00
We will soon be in a position to start our TS2 checks to enable us to
bring it online. The machine will be intermittent whilst we do this
operation.
20 February 2026 07:07
Beam ON
20 February 2026 06:30
After contacting equipment experts we replaced the failed steering magnet
power supply. However, this was not successful and so equipment experts
are now on their way to site to investigate further.
Update @ 07:30.
20 February 2026 05:50
BEAM OFF - There is an issue with one of the synchrotron steering magnet
power supplies and we have gained access to the inner synch to investigate
further.
Update @ 06:30.
19 February 2026 10:39
Machine update:-
Some target 2 beam intensity has been displayed on the beam intensity
page. This is due to calibration testing and setup of the intensity
monitors. No beam is being transported to target 2, so please ignore any
displays whilst setup is continuing.
Target 2 repairs will be continually updated.
17 February 2026 21:21
Machine update:-
Beam restored to Target 1.
17 February 2026 20:30
Machine update:-
Experts are on site trying to rectify the fault. The fault has been
identified with a power supply and is being changed at the moment. Update
21:00hrs.
17 February 2026 19:05
Machine update:-
We have problems with the extract kickers. Experts have been contacted and
are on route to help with repairs. Update 20:30hrs
17 February 2026 13:20
The following Beamlines are also cleared for beam:
RIKEN PORT 2
RIKEN PORT 3
RIKEN PORT 4
RIKEN KICKER
MUON HIFI
MUON MuSR
MUON EMU
MUON BENDING MAGNET
N6 IRIS/OSIRIS MAIN
N6A IRIS
17 February 2026 10:44
Attention all Target Station 1 users.
The following beamlines are cleared for beam:
N1 SANDALS
N2 ALFRED
N3 SURF
N4 CRISP
N5 LOQ
N7 POLARIS
N8 TOSCA
N8A INES
N9 EMMA
S1 MAPS
S2 VESUVIO
S3 SXD
S4 MERLIN
S6 MARI
S7 GEM
S8P ENGIN-X
S8X ENGIN-X
S9 PEARL
17 February 2026 08:30
User Run 2025/5 has started.
20 December 2025 07:55
Beam to SYNCHROTRON ROOM BEAM DUMP has finished. Access to the R6 area can
continue as normal.
19 December 2025 20:01
At 21:00 Beam will be set to SYNCHROTRON ROOM BEAM DUMP. Please do not
enter R6 north end via chalk pits, R2 lift to tunnel 3 or R6 bat cave.
Sign have been placed in these areas.
19 December 2025 08:32
User run ended at 08:30.
19 December 2025 00:08
Update:
The issue with the synch RF liquid resistor flows has now been resolved.
BEAM ON
18 December 2025 21:42
Expert is on route to investigate. Although he has said fix could take
several hours.
Update @ 00:30
18 December 2025 21:28
There is an issue with one of the synch RF liquid resistor flows.
Currently contacting an expert. Update @ 22:00
18 December 2025 15:00
TS2 H2 USERS - The H2 moderator is continuing to cool down and should be
at an operational temperature level in the next few hours. Please continue
to monitor the temperature and an update will be issued once the
temperature reaches an operational level.
18 December 2025 13:08
TS2 USERS - Beam has been restored to TS2. However, the H2 moderator
temperature is still high but should continue to cool down over the next
few hours.
Update @ 15:00
18 December 2025 11:12
Beam has been restored to TS1.
TS2 USERS - As a result of the mains dip the H2 compressor tripped off
which has caused the H2 moderator to VETO and warm up. Equipment experts
are working to get the system running again before they can cool down the
moderator.
Update @ 13:00.
18 December 2025 10:33
BEAM OFF - We have suffered a mains dip that has affected a large amount
of operational equipment. We are in the process of getting equipment back
on and run up but this will take some time.
Update @ 11:30.
18 December 2025 10:05
Beam ON
18 December 2025 09:17
BEAM OFF - We are experiencing an issue with 1RF2 liquid resistor circuit.
Equipment experts are in the inner synchrotron working on the fault.
Update @ 10:00.
17 December 2025 10:44
Beam running back up.
17 December 2025 10:04
Beam currently off while investigating a diagnostic issue. Update @ 10:30
17 December 2025 06:14
Update:
The issue with the collectors has now been resolved.
BEAM ON.
17 December 2025 05:32
Update:
We are currently investigating an issue with the collectors that looks as
though it has been causing the increased beam loss in the Synchrotron.
Update at 06:30
17 December 2025 04:09
BEAM OFF - We are experiencing high beam loss in the Synchrotron. A machine
expert is on route to site to assist. Update at 05:30
16 December 2025 17:32
Beam ON
16 December 2025 17:00
BEAM OFF - We are experiencing high beam loss in the Synchrotron. Machine
experts are currently investigating the cause of this.
Update @ 18:00
16 December 2025 06:39
Target 2 users - Compressor purging and cryo-system cool down has been
completed quicker than expected.
BEAM ON to both targets.
15 December 2025 12:36
Target 2 user. Compressor purging and cryo-system cool down is expected to
take approx 2 days. If all goes well Beam should return Wednesday morning.
15 December 2025 10:44
Target 2 users. The valve in the compressor that required replacing has
been changed. Systems are currently being purged, the compressor should be
ready for a start within the hour. Update @ 12:00
14 December 2025 19:03
Experts believe they have worked out the issue with Target 2 compressor, a
failed valve. A replacement valve is currently unavailable this evening,
they are hoping a replacement can be sourced early tomorrow morning. No
ISIS operating schedule
Find out about our upcoming operating schedule. Cycles begin/end at 08:30 on the date specified.
| Cycle | Start | End | Target station(s) | Maintenance day(s) | Comments |
|---|---|---|---|---|---|
| 2025/5 | 17 Feb 2026 | 27 Mar 2026 | TS1 & TS2 | 11 Mar 2026 | - |
| 2026/1 | 05 May 2026 | 12 Jun 2026 | TS1 & TS2 | 20 May 2026 | - |
| 2026/2 | - | - | - | - | Cycle cancelled |
| 2026/3 | 15 Sep 2026 | 23 Oct 2026 | TS1 & TS2 | 30 Sep 2026 | - |
| 2026/4 | 17 Nov 2026 | 18 Dec 2026 | TS1 & TS2 | 02 Dec 2026 | - |
| 2026/5 | 02 Feb 2027 | 12 Mar 2027 | TS1 & TS2 | TBC | - |
ISIS operating schedule - future years
From April 2026 to March 2027, we expect to run four cycles which will be allocated in proposal rounds 26-1 (deadline October 2025) and 26-2 (deadline April 2026).
We are currently anticipating having an extended shutdown for accelerator and beamline work from March to early September 2027. This would mean no proposal round in October 2026, and a round in April 2027 to allocate time from September 2027 to March 2028.
Machine physics schedule
Accelerator physics experiments are performed before and after each user run. The latest schedule can be seen by clicking this text.
| Date | Shift | Mode | Rep Rate | Task | Contacts | Notes |
| 3/2/26 Tuesday |
00:00–08:30 | 5 | — | – | – | – |
| 3/2/26 Tuesday |
08:30–14:00 | 5 | MS/32 | Initial Injector Setup | HVC, SK | Intermittent beam to IHBS2 |
| 3/2/26 Tuesday |
14:00-16:30 | 5 | MS/32 | Commissioning IPM 11, 12, 13 | SAF, SK | Intermittent beam to IHBS2 |
| 3/2/26 Tuesday |
16:30–00:00 | 5 | MS/32 | Initial Injector Setup and HEDS Model Measurements | BK | Intermittent beam to IHBS2 |
| 4/2/26 Wednesday |
00:00–08:30 | 5 | – | – | – | – |
| 4/2/26 Wednesday |
08:30–12:30 | 5 | MS/32 | Orbit Response Matrix Measurements | REW, EA | Storage Ring Mode HEDS Chopper Required |
| 4/2/26 Wednesday |
12:30-14:30 | 5 | MS/32 | Extract Septum Scans in Bunched Storage Ring Mode | REW | Bunched Storage Ring Mode |
| 4/2/26 Wednesday |
14:30-18:30 | 5 | MS/32 | AC Dipole Testing | CJ, DPdB | Storage Ring Mode |
| 4/2/26 Wednesday |
18:30–00:00 | 5 | MS/32 | Impedance Measurements | DPdB | Storage Ring Mode |
| 5/2/26 Thursday |
00:00–08:30 | 5 | MS/32 | Run up MMPS DC and AC, turn on RF | Crew | – |
| 5/2/26 Thursday |
08:30–16:30 | 5 | MS/32 | Initial Synchrotron Setup | BK, SK | – |
| 5/2/26 Thursday |
16:30–00:00 | 5 | MS/32 | Synchrotron Setup | CMW, EA | – |
| 6/2/26 Friday |
00:00–08:30 | 5 | MS/32 | – | – | – |
| 6/2/26 Friday |
08:30–16:30 | 5 | MS | Synchrotron Setup & LEBT optimisation tests | REW, HR, RB | – |
| 6/2/26 Friday |
16:30-18:30 | 5 | MS | SCPM Tests | DAb | – |
| 6/2/26 Friday |
18:30–00:00 | 5 | MS | Synchrotron Setup & EPB Alignments | HVC | – |
| 7/2/26 Saturday |
00:00–08:30 | 5 | MS | – | – | – |
| 7/2/26 Saturday |
08:30-16:30 | 5 | MS | Synchrotron Setup | REW, SK | – |
| 7/2/26 Saturday |
16:30–00:00 | 5 | MS | Synchrotron Setup | DJA | – |
| 8/2/26 Sunday |
00:00–08:30 | 5 | MS | – | – | – |
| 8/2/26 Sunday |
08:30–16:30 | 5 | MS | Synchrotron Setup | HVC, EA | – |
| 8/2/26 Sunday |
16:30–00:00 | 5 | MS | Synchrotron Setup | BK | – |
| 9/2/26 Monday |
00:00–08:30 | 5 | MS | – | – | – |
| 9/2/26 Monday |
08:30–12:30 | 5 | MS | Synchrotron Setup | HVC | – |
| 9/2/26 Monday |
12:30-16:30 | 5 | MS | 1RF Bucket Shape Tests | AS, DJA, (HVC) | – |
| 9/2/26 Monday |
16:30–00:00 | 5 | MS | RCS Chopped Beam dQ/dI Measurements | HR | HEDS Chopper Required |
| 10/2/26 Tuesday |
00:00–08:30 | 5 | MS | – | – | – |
| 10/2/26 Tuesday |
08:30–16:30 | 5 | MS | RF Setup & Aperture Measurements | AS, ISKG, HVC | – |
| 10/2/26 Tuesday |
16:30–00:00 | 5 | MS | Synchrotron Setup | DJA | – |
| 11/2/26 Wednesday |
00:00–08:30 | 5 | MS | – | – | – |
| 11/2/26 Wednesday |
08:30–14:00 | 5 | MS | Synchrotron Setup | HVC, HR | – |
| 11/2/26 Wednesday |
14:00-16:30 | 5 | MS | Bunch Compression Setup (Theta Method) | BK, AS | – |
| 11/2/26 Wednesday |
16:30–00:00 | 5 | MS | Synchrotron Setup | CMW | – |
| 12/2/26 Thursday |
00:00–08:30 | 5 | MS | – | – | – |
| 08:30–14:00 | 5 | MS | Synchrotron Setup | HVC, SK | – | |
| 12/2/26 Thursday |
14:00-16:30 | 5 | MS | Diagnostics Measurements: Intensity monitoring, TS2 BLMs TS-1 Timing Tests |
JFG CS |
– |
| 12/2/26 Thursday |
16:30–00:00 | 5 | MS | Synchrotron Setup | BK, EA | – |
| 13/2/26 Friday |
00:00–08:30 | 5 | MS | – | – | – |
| 13/2/26 Friday |
08:30–16:30 | 5 | MS | Synchrotron Setup, incl. Halo steering tests | REW, DK, IF | – |
| 13/2/26 Friday |
16:30-18:30 | 5 | MS | LEBT Space Charge Compensation Tests | EF, OT | – |
| 13/2/26 Friday |
18:30–00:00 | 5 | MS | Synchrotron Setup & Chopped Injection Scan | HVC | HEDS Chopper Required |
| 14/2/26 Saturday |
00:00–08:30 | 5 | MS | – | – | |
| 14/2/26 Saturday |
08:30–16:30 | 5 | MS | Synchrotron Setup | BK | – |
| 14/2/26 Saturday |
16:30–00:00 | 5 | MS | Synchrotron Setup | DJA | – |
| 15/2/26 Sunday | 00:00–08:30 | 5 | MS | – | – | – |
| 15/2/26 Sunday |
08:30–16:30 | 5 | MS | Synchrotron Setup | HVC | – |
| 15/2/26 Sunday |
16:30–00:00 | 5 | MS | Beamline Surveys | HP Section | – |
Previous cycles
| Cycle | Start | End | Target station(s) | Maintenance day(s) | Comments |
|---|---|---|---|---|---|
| 2025/4 | 18 Nov 2025 | 19 Dec 2025 | TS1 & TS2 | 03 Dec 2025 | |
| 2025/3 | 09 Sep 2025 | 17 Oct 2025 | TS1 & TS2 | 24 Sep 2025 | |
| Cycle cancelled | |||||
| 2025/1 | 29 Apr 2025 | 06 Jun 2025 | TS1 & TS2 | 14 May 2025 | |
| 2024/5 | 18 Feb 2025 | 28 Mar 2025 | TS1 & TS2 | 05 Mar 2024 | |
| 2024/4 | 19 Nov 2024 | 20 Dec 2024 | TS1 & TS2 | 04 Dec 2024 | |
| 2024/3 | 17 Sep 2024 | 20 Dec 2024 | TS1 & TS2 | 04 Dec 2024 | |
| 2024/2 | 09 Jul 2024 | 28 Mar 2025 | TS1 & TS2 | 05 Mar 2024 | |
| 2024/1 | 23 Apr 2024 | 31 May 2024 | TS1 & TS2 | 08 May 2024 | |
| 2023/5 | 13 Feb 2024 | 22 Mar 2024 | TS1 & TS2 | 28 Feb 2024 | |
| 2023/4 | 14 Nov 2023 | 19 Dec 2023 | TS1 & TS2 | 29 Nov 2023 | |
| 2023/3 | 19 Sep 2023 | 20 Oct 2023 | TS1 & TS2 | 04 Oct 2023 | |
| 2023/2 | 27 Jun 2023 | 04 Aug 2023 | TS1 & TS2 | 12 Jul 2023 | |
| 2023/1 | 25 Apr 2023 | 26 May 2023 | TS2 Only (TS1 Commissioning) | 10 May 2023 |